# Explainable Multi-Modal Data Exploration in Natural Language via LLM Agent

### v1

@@ 4.1.1. Datasets
Dataset 1: Artwork. We use the artwork dataset introduced by (urbanB24, ). This dataset contains information about paintings in tabular form as well as an image collection containing 100 images of the artworks. This data is taken from Wikipedia. The tabular data contains metadata information about paintings such as title, inception, movement, etc. as well as a reference to the respective paintings. A typical example question from this dataset is Plot the number of paintings depicting war for each century (as previously shown in Figure 3). In addition to the 24 existing questions in the artwork dataset, we propose six new questions aimed at evaluating parallel task planning and execution, facilitating a comparison between the characteristics of the two architectures. These six questions incorporate both single and multiple modalities. Moreover, four of the six questions require responses in various formats: two questions demand two plots, and two questions involve a combination of plotting and showing the results in a specific data structure, i.e. either as a tabular format or as a JSON format. The final test dataset contains 30 natural language questions derived from the original 24 in the artwork dataset. These include 8 queries seeking a single result value, 11 requiring structured data as output, and 11 requesting a plot. Of these, 18 queries involve multi-modal data, while the remaining 12 are based exclusively on relational data.
We have chosen this dataset to directly compare our system with CAESURA (urbanB24, ), one of the state-of-the-art systems for multi-modal data exploration in natural language.
@@
Dataset 2: Electronic Health Records (EHR). We also utilized the EHRXQA (bae2024ehrxqa, ) dataset, a multi-modal question answering dataset that integrates structured electronic health records (EHRs) with chest X-ray images. This dataset consists of 18 tables and 432 images, and specifically requiring cross-modal reasoning. The questions of EHRXQA are categorized based on their scope in terms of modality and patient relevance. For modality-based categorization, questions were classified into three types: Table-related, image-related, and table-image-related, based on the data modality required. The patient-based categorization classified questions based on their relevance to a single patient, a group of patients, or none (i.e., unrelated to specific patients). We have chosen this dataset since it was used to evaluate NeuralSQL, another state-of-the-art system for multi-modal data exploration. To manage the cost of an API call, we extracted randomly 100 questions. The selection process was guided by three predefined categories within the test set of the EHRXQA dataset: Image Single-1, Image Single-2, and Image+Table Single.
Here are examples from each category, taken from the original paper (bae2024ehrxqa, ). All questions in these categories require multi-modal data exploration for the reasoning process.
  • Image Single-1: Given the last study of patient 15439, which anatomical finding is associated with the right lower lung zone, pneumothorax or vascular redistribution?
  • Image Single-2: Enumerate all newly detected diseases in the last study of patient 19290 in 2103 compared to the previous study.
  • Image+Table Single: Did a chest X-ray study for patient 15110 reveal any anatomical findings within 2 months after the prescription of hydralazine since 2021?
4.1.2. Baseline Systems and Setup
@@
We compare XMODE to the baseline implementations of CAESURA (urbanB24, ) and NeuralSQL (bae2024ehrxqa, ) - two important state-of-the-art systems for multi-modal data exploration.
@@
CAESURA supports natural language queries over a multi-modal data lake leveraging BLIP-2 (li2023blip, ) for visual question answering and BART (lewis2020bart, ) for text question answering. We reproduced the results of CAESURA on the Artwork dataset using GPT4o. To compare to our system, we utilize GPT4o as an LLM and the same model for visual question answering (i.e., BLIP-2) in XMODE.
@@
In NeuralSQL, an LLM is integrated with an external visual question answering system, M3AE model  (10.1007/978-3-031-16443-9_65, ), to handle multi-modal questions over a structured database with images by translating a user question to SQL in one step. To ensure that we used the optimal hyperparameter settings and prompt structure, we contacted the authors of EHRXQA (bae2024ehrxqa, ), who provided the results of their experiment for NeuralSQL using GPT-4o on 100 randomly selected questions.
For XMODE, we employ the M3AE model with task-specific fine-tuned weights, provided by  (bae2024ehrxqa, ), for the image analysis task. The customized M3AE model is encapsulated as a web service and is deployed on the same computing node described in Section 4.1.4
4.1.3. Evaluation Metrics
To evaluate XMODE against state-of-the-art systems, we use the following metrics:
  • Accuracy: Measures the accuracy (i.e., exact match) of the generated result set compared with the gold standard result set or with the human expert.
  • Steps: Number of steps required by the respective system to come up with the final result. These steps include reasoning, planning, re-planning etc.
  • Tokens: Number of tokens used for prompt engineering.
  • Latency: End-to-end execution time for a system to come up with the final result.
  • API costs: Costs for calling the LLM, e.g. for GPT4o.
We apply the above-mentioned metrics under various question and system categories:
  • Modality: Questions can either be of single modality, i.e. querying only relational data or image data, or of multiple modality, i.e. querying both relational and image data.
  • Output Type: The output type of a question can either be a single value, e.g. true or false, a data structure, e.g. in tabular or JSON format, a plot, or a combination of plots and data structures.
  • Workflow: The generated workflow plan can either be sequential or parallel.
Finally, we evaluate if a system generates a correct (multi-modal) query plan (i.e. generated plan), and if it supports re-planning.
4.1.4. Hardware Setup
@@
We conduct the following experiments using a CUDA-accelerated computational node on an OpenStack virtual host. This node is equipped with a 16-core CPU, 16 GB of main memory, and 240 GB of SSD storage. Additionally, it features an NVIDIA T4 GPU with 16 GB of dedicated graphics memory.
4.2. Results on the Artwork Dataset
We first evaluate the results on the artwork dataset and afterward on the EHR dataset.
4.2.1. Performance Results
System Category Accuracy Steps Tokens Latency [s] API Cost [USD] Generated Plan Re-planning
CAESURA Modality Single (15) 60.00% 152 214,014 973.28 1.33 80% No
Multiple (15) 6.67% 164 268,918 4,847.95 1.65
Output Type Single Value (8) 37.50% 88 135,077 1,047.24 0.82
Data Structure (10) 50.00% 116 183,454 2,683.03 1.14
Plot (8) 25.00% 79 112,732 1,856.66 0.69
few-shot ($n=4$) Plot-Plot (2) 0% 16 21,508 108.87 0.14
in planning Plot-Data Structure (2) 0% 17 30,161 125.42 0.19
Workflow Sequential (24) 41.67% 261 399,045 5,330.12 2.45
Parallel (6) 0% 55 83,887 491.11 0.52
Overall (30) 33.33% 316 482,932 5,821.23 2.98
XMODE Modality Single (15) 100.00% 96 159,212 525.09 0.61
Multiple (15) 26.67% 107 326,400 2,515.03 1.49 100% Yes
Output Type Single Value (8) 50.00% 56 71,575 494.78 0.39
Data Structure (10) 50.00% 67 223,528 1,330.40 0.89
Plot (8) 75.00% 52 118,431 798.97 0.48
zero-shot Plot-Plot (2) 100.00% 14 50,108 308.92 0.22
Plot-Data Structure (2) 100.00% 14 21,970 107.05 0.10
Workflow Sequential (24) 62.50% 163 338,766 2,131.11 1.51
Parallel (6) 66.67% 40 146,846 909.01 0.59
Overall (30) 63.33% 203 485,612 3,040.12 2.10
Table 1. Performance metrics of Caesura (urban2023caesura, ) and XMODE on the artwork dataset.
Table 1 presents a comparison of XMODE and CAESURA on the artwork dataset across various aspects. The performance metrics for each aspect were determined through a manual evaluation conducted by our team of four researchers. The comparison between XMODE and CAESURA reveals notable differences in their performance across various aspects of the artwork dataset. Starting with the metric accuracy for evaluating queries of different modalities: XMODE outperforms CAESURA in both single- and multi-modality questions. For single cases, XMODE achieves an output accuracy of 100.00%, while CAESURA falls behind at 60.00%. In the more challenging multiple-modality scenarios, XMODE demonstrates a significant edge with 26.67% accuracy, compared to CAESURA’s much lower 6.67%.
The accuracy based on the output types shows that for single-value outputs CAESURA achieves an accuracy of 37.5%, while XMODE yields 50%. XMODE’s edge is even more evident for complex tasks where the output is plot, plot-plot, or plot-data structure. Here XMODE reaches an accuracy of 75%, 100%, and 100%, respectively, significantly surpassing CAESURA’s performance of 25%, 0%, and 0%, respectively.
Moreover, CAESURA requires sequential reasoning and acting for each natural language question which can result in high latency, cost, and sometimes inaccurate behavior. XMODE identifies dependencies between tasks during workflow planning and thus enables concurrent and parallel task execution. On a set of six new questions requiring parallel task planning and execution, CAESURA fails entirely, while XMODE successfully generates proper plans and achieves 66.67% accuracy.
Overall, XMODE emerges as the stronger system, with an overall output accuracy of 63.33%, compared to CAESURA’s 33.33%. XMODE distinguishes itself with its ability to provide better explanations, support re-planning, and concurrency—features absent in CAESURA.
From an efficiency perspective, XMODE demonstrates significant advantages over CAESURE in several areas. It requires fewer steps (203 vs. 316) and achieves significantly lower latency (3,040.12 ms vs. 5,821.23 ms) demonstrating a faster response time. Finally, XMODE also incurs a lower API cost (2.10) compared to CAESURA’s 2.98, indicating that XMODE is more cost-effective in terms of API usage.
In summary, the experiment results using the artwork benchmark showed that XMODE consistently outperforms CAESURA in accuracy, efficiency, and feature support, demonstrating its robustness across a variety of tasks on the artwork dataset. Its ability to handle complex outputs, provide explanations, and adapt through replanning positions it as the better choice in this benchmark.
4.2.2. Optimizations of XMODE Explained with Examples
To better demonstrate advantages of XMODE, we provide several examples (see Figures 3 and 5) across three key aspects: explanations, smart replanning, and parallel planning. The following examples provide a detailed illustration of these three aspects.
{mdframed}
[hidealllines=true,backgroundcolor=cyan!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-1pt,rightmargin=-1pt]
Figure 5. Optimization of XMODE: Smart replanning.
3) Decision Making: When the tasks are completed, the model reflects on its work and provides a final output based on its thought as Summary:"The number of paintings depicting war has been plotted for the 16th and 18th centuries.", "Details": "The analysis identified 1 painting from the 16th century and 2 paintings from the 18th century that depict war. The plot visualizes these findings. [..]”. Throughout the workflow, the model demonstrates a commitment to transparency.
At every stage, XMODE provides reasoning to justify its actions, from choosing SQL for retrieval to selecting a bar chart for visualization. Intermediate outputs, like the dataset of war paintings and the Python plotting code, are made visible, ensuring the user can trace the steps taken. The decision making phase wraps up the process by summarizing findings, clarifying the approach, and sharing the final visual result. This shows that XMODE not only answers the query effectively but also ensures its steps are understandable, logical, and well-documented, building trust in its analysis.
Figure 6. Optimization of XMODE: Parallel planning.
{mdframed}
[hidealllines=true,backgroundcolor=cyan!20,innerleftmargin=2pt,innerrightmargin=3pt,leftmargin=-1pt,rightmargin=-1pt] Example 2 - Smart Replanning: What is depicted on the oldest Renaissance painting in the database? (see Figure 5). Contrary to the previous example, XMODE here involves smart replanning - a major optimization technique of XMODE. The main idea is to dynamically adapt the planning in case some tasks of the workflow fail or do not produce any results. Here’s a breakdown of each step:
1) Planning & Expert Model Allocation: XMODE outputs the initial workflow plan that has 2 tasks. The first task involves retrieving the image path and the year of the oldest Renaissance painting in the database using a ”text2SQL” expert model. It also involves an ”image_analysis” expert model in the second task, which aims to determine what is depicted in the image.
2) Execution and Self-Debugging: XMODE takes the information about the planned workflow as well as task dependencies and puts it into action. In Task 1, it comes with a reasoning statement to generate the SQL query as: SELECT img_path, strftime(’%Y’, inception) AS year FROM paintings WHERE movement = ’Renaissance’ ORDER BY inception ASC LIMIT 1. Then it executes the query over the Artwork database and retrieves the specific image path and year for the oldest Renaissance painting as [’img_path’: ’images/img_0.jpg’, ’year’: ’1438’]. This allows the model to access the actual painting data in the subsequent task.
In Task 2, XMODE utilizes the ”image_analysis” expert model (i.e. visual question answering based on BLIP) to examine the contents of img_0.jpg to answer the question: What is depicted in the image? The output of this task is transferred as a final result to the decision making component. At this point, the model’s ”thought” process in this component becomes evident. It reasons that while it knows that img_0.jpg is a painting, the details about what is depicted in the painting have not been provided. Therefore, the model decides to not provide a final answer to the user and does replanning.
The replanning capability is a crucial aspect of the XMODE’s approach. Rather than blindly accepting the final answer which does not produce a satisfiable or correct result, the model recognizes the need to replan and calls the ”image_analysis” module again. Since the model already knows which image in the database contains the oldest Renaissance painting, it smartly plans the ”image_analysis” task as Task 3, by reformulating the question as What is specifically depicted in the painting? XMODE then executes the task, and receives the more concrete answer ”umbrellas”.
Moving forward, the decision making component confirms the details about the painting. Here, it verifies that the information it has gathered so far aligns with the natural language question and makes sense as a comprehensive understanding of the oldest Renaissance painting. The key aspect is the model’s ability to replan effectively and to strategically leverage the available information to avoid repeating tasks. {mdframed}[hidealllines=true,backgroundcolor=cyan!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-1pt,rightmargin=-1pt] Example 3 - Parallel Planning: In the Renaissance, find the total number of paintings depicting war and the number of paintings depicting swords (see Figure 6). The figure illustrates how XMODE process a complex query about Renaissance paintings, focusing on identifying how many paintings depict war and how many depict swords. The pipeline is structured to combine parallel task execution with step-by-step explanations, ensuring clarity and efficiency throughout the process.
The process begins in the Planning & Expert Model Allocation, where the model breaks down the user’s query into distinct subtasks. These subtasks are assigned to specialized modules: Task 1 ”text2SQL”: This task retrieves image paths and relevant metadata for Renaissance paintings from a database using a SQL query. Task 2 ”image_analysis”: This task examines whether each painting depicts war. Task 3 ”image_analysis”: Simultaneously, another module analyzes whether each painting depicts a sword. Task 4 ”data_preparation”: This task consolidates the results from Task 2 and Task 3 to count and summarize the paintings.
At this point, the model showcases its parallel planning capability. Tasks 2 and 3 are performed concurrently: For Task 2, the system uses image analysis to determine if each painting depicts war. For Task 3, a similar image analysis process identifies paintings that depict swords. Running these tasks in parallel significantly speeds up the workflow, as they operate independently of each other. Once the image analysis tasks are complete, the model transitions to Task 4, where it aggregates the results. The reasoning here details how the system compiles two lists - one for paintings depicting war and one for those depicting swords. Afterwards, XMODE counts the entries in each list. The final results are prepared for the decision making module.
In the decision making phase, the model reflects on its findings. It confirms that sufficient data was processed to answer the query and provides a summary: "There is 1 painting depicting war and 38 paintings depicting swords.
XMODE offers details, explaining how the analysis was conducted and highlighting the disparity between the two categories of paintings. The system further provides an explanation of its methodology, emphasizing how it worked systematically to answer the query. This demonstrates XMODE’s ability to manage tasks efficiently through parallel execution and to ensure transparency through reasoned explanations at every step. By combining these capabilities, the system provides a clear, accurate, and well-supported response to the user’s query.
It is important to note that we did not compare XMODE to NeuralSQL on ArtWork’s questions, as such a comparison would be unfair due to NeuralSQL’s inability to support plotting.
4.3. Results on the EHRXQA Dataset
In this section, we evaluate the performance of NeuralSQL and XMODE on the EHRXQA dataset. This comparison excludes metrics like steps, tokens, and latency because evaluating XMODE’s performance on these aspects against NeuralSQL is not meaningful. NeuralSQL generates the final answer in a single step without providing a plan or intermediate steps, whereas our approach focuses on decomposing natural language questions, planning workflows, and responding transparently.
We also exclude CAESURA from the EHRXQA experiments. While CAESURA is intended to be a general-purpose multimodal system, it processes the relational database through multiple steps, examining each table and relationship sequentially. This limitation introduces significant overhead when handling the complex data schema of the EHRXQA dataset (there are 18 tables) during the discovery phase. Consequently, reproduing CAESURA on EHRXQA questions fails to perform inferences at the early stages of the planning phase, ultimately terminating after exceeding the maximum number of allowed attempts.

### v2

@@ 1. Dataset 1:
    In addition to the 24 existing questions in the ArtWork dataset, we propose six new questions aimed at evaluating parallel task planning and execution, facilitating a comparison between the characteristics of the two architectures. These six questions incorporate both single and multiple modalities. Moreover, four of the six questions require responses in various formats: two questions demand two plots, and two questions involve a combination of plotting and showing the results in a specific data structure, i.e. either as a tabular format or as a JSON format. The final test dataset contains 30 natural language questions derived from the original 24 in the ArtWork dataset. These include 8 queries seeking a single result value, 11 requiring structured data as output, and 11 requesting a plot. Of these, 18 queries involve multi-modal data, while the remaining 12 are based exclusively on relational data. We have chosen this dataset to directly compare our system with CAESURA (Urban and Binnig, 2024), one of the state-of-the-art systems for multi-modal data exploration in natural language.
    Figure 2: M2EX system architecture in EHRXQA  (Bae et al., 2024) with an example of processing a multi-modal query. The query is automatically decomposed into various components which can be inspected by the user for explainability.
  2. Dataset 2:
    RotoWire. This dataset is also utilized by Urban and Binnig (2024) and consists of one relational database and 100 randomly selected textual reports about NBA games, including metadata, key statistics of individual players, and team performance metrics. A typical example question from this dataset is Plot the highest number of three-pointers made by players from each nationality. The test dataset comprises 12 natural language questions, evenly divided into 6 single-modal and 6 multi-modal queries. Regarding output format, 3 questions require a single value as a response, 5 involve structured data outputs, and 4 necessitate visualization through plots.
  3. Dataset 3:
@@
    Electronic Health Records (EHR). We also utilized the EHRXQA (Bae et al., 2024) dataset, a multi-modal question answering dataset that integrates structured electronic health records (EHRs) with chest X-ray images. This dataset consists of 18 tables and 432 images, and specifically requires cross-modal reasoning. The questions of EHRXQA are categorized based on their scope in terms of modality and patient relevance. For modality-based categorization, questions were classified into three types: Table-related, image-related, and table-image-related, based on the data modality required. The patient-based categorization classified questions based on their relevance to a single patient, a group of patients, or none (i.e., unrelated to specific patients). We have chosen this dataset since it was used to evaluate NeuralSQL, another state-of-the-art system for multi-modal data exploration. To manage the cost of an API call, we extracted 100 questions randomly. The selection process was guided by three predefined categories within the test set of the EHRXQA dataset: Image Single-1, Image Single-2, and Image+Table Single (for details, please look at Bae et al. (2024)).
Several considerations influenced our decision to work with reduced versions of these datasets: Demonstrating ViabilityThe reduced dataset size demonstrates M2EX’s viability across diverse multi-modal datasets with ground truth, proving its ability to handle complex queries in a controlled setting. Complexity of Building DatasetsConstructing large-scale multi-modal datasets with precise ground truth is a complex, manual process, which limits the scaling-up within the study’s scope. Cost ConsiderationsThe cost of API calls to the LLM powering M2EX necessitates a balance between dataset size and experimental feasibility, ensuring thorough evaluation within practical constraints.
4.2 Baseline Systems and Setup
@@
We compare M2EX to the baseline implementations of CAESURA (Urban and Binnig, 2024) and NeuralSQL (Bae et al., 2024) - two important state-of-the-art systems for multi-modal data exploration.
@@
CAESURA supports natural language queries over a multi-modal data lake, leveraging BLIP-2 (Li et al., 2023b) for visual question answering and a fine-tuned BART (Lewis et al., 2020) for text question answering. We reproduced the results of CAESURA on the ArtWork and RotoWire datasets using GPT-4o for planning, data processing, and plot generation while adopting the other tool models as proposed in CAESURA (Urban and Binnig, 2024). For comparison with our system, we use GPT-4o as the LLM for both planning and text analysis on RotoWire. On ArtWork, we employ GPT-4o as the planner and retain the same model for visual question answering (i.e., BLIP-2) in M2EX.
@@
In NeuralSQL, an LLM is integrated with an external visual question answering system, M3AE model (Chen et al., 2022), to handle multi-modal questions over a structured database with images by translating a user question to SQL in one step. To ensure that we used the optimal hyperparameter settings and prompt structure, we contacted the authors of EHRXQA (Bae et al., 2024), who provided the results of their experiment for NeuralSQL using GPT-4o on 100 randomly selected questions.
@@
For M2EX, we employ the M3AE model with task-specific fine-tuned weights, provided by  (Bae et al., 2024), for the image analysis task. The customized M3AE model is encapsulated as a web service and deployed on the same computing node as our experiments. We conduct the experiments using a CUDA-accelerated computational node on an OpenStack virtual host. This node is equipped with a 16-core CPU, 16 GB of main memory, and 240 GB of SSD storage. Additionally, it features an NVIDIA T4 GPU with 16 GB of dedicated graphics memory. A complete mapping of subtasks to expert models/tools and prompt types is provided in Table 4 (Appendix C).
4.3 Evaluation Metrics
To evaluate M2EX against state-of-the-art systems, we use the following metrics: (i) Accuracy: Measures the accuracy (i.e., exact match) of the generated result set compared with the gold standard result set or with the human expert. (ii) Steps: Number of steps required by the respective system to come up with the final result. These steps include reasoning, planning, re-planning, etc. (iii) Tokens: Number of tokens used for prompt engineering. (iv) Latency: End-to-end execution time for a system to come up with the final result. (v) API costs: Costs for calling the LLM, e.g. for GPT4o.
We apply the above-mentioned metrics under various questions and system categories:
(i) Modality: Questions can either be of single modality, i.e., querying only relational data or image data, or of multiple modalities, i.e., querying both relational and image data. (ii) Output Type: The output type of a question can either be a single value, e.g., true or false, a data structure, e.g., in tabular or JSON format, a plot, or a combination of plots and data structures. (iii) Workflow: The generated workflow plan can either be sequential or parallel.
Finally, we evaluate if a system generates a correct (multi-modal) query plan (i.e., generated plan), and if it supports re-planning.
4.4 Results on the Benchmark Datasets
System Category (# in ArtWork||# in RotoWire) ArtWork RotoWire Re-planning
Accuracy Steps Tokens Latency [s] Cost [\$] Gen. Plan Accuracy Steps Tokens Latency [s] Cost [\$] Gen. Plan
CAESURA Modality Single ($15|6$) 60.00% 152 214,014 973.28 1.33 80% 50.00% 79 100,277 500.52 0.65 91.67% No
Multiple ($15|6$) 6.67% 164 268,918 4,847.95 1.65 0.00% 78 133,230 959.17 0.85
Output Single Value ($8|3$) 37.50% 88 135,077 1,047.24 0.82 66.67% 32 45,145 287.55 0.29
Data Structure ($10|5$) 50.00% 116 183,454 2,683.03 1.14 20.00% 69 104,345 659.37 0.68
Plot ($8|4$) 25.00% 79 112,732 1,856.66 0.69 0.00% 56 84,017 512.77 0.53
few-shot ($4$) Type Plot-Plot ($2|0$) 0% 16 21,508 108.87 0.14
in planning Plot-Data Structure ($2|0$) 0% 17 30,161 125.42 0.19
Workflow Sequential ($24|12$) 41.67% 261 399,045 5,330.12 2.45 25.00% 157 233,507 1,459.69 1.50
Parallel ($6|0$) 0% 55 83,887 491.11 0.52
Overall ($30|12$) 33.33% 316 482,932 5,821.23 2.98 25.00% 157 233,507 1,459.69 1.50
M2EX Modality Single ($15|6$) 100.00% 96 159,212 525.09 0.61 100.00% 34 89,810 524.06 0.40
Multiple ($15|6$) 26.67% 107 326,400 2,515.03 1.49 100% 33.33% 42 952,386 3,235.96 3.22 100% Yes
Output Single Value ($8|3$) 50.00% 56 71,575 494.78 0.39 100.00% 16 108,520 499.70 0.40
Data Structure ($10|5$) 50.00% 67 223,528 1,330.40 0.89 40.00% 27 410,698 2,120.15 1.57
Plot ($8|4$) 75.00% 52 118,431 798.97 0.48 75.00% 33 522,987 1,140.17 1.65
zero-shot Type Plot-Plot ($2|0$) 100.00% 14 50,108 308.92 0.22
Plot-Data Structure ($2|0$) 100.00% 14 21,970 107.05 0.10
Workflow Sequential ($24|12$) 62.50% 163 338,766 2,131.11 1.51 66.67% 76 1,042,196 3,760.02 3.62
Parallel ($6|0$) 66.67% 40 146,846 909.01 0.59
Overall ($30|12$) 63.33% 203 485,612 3,040.12 2.10 66.67% 76 1,042,196 3,760.02 3.62
Table 1: Performance metrics of Caesura (Urban and Binnig, 2024) and M2EX on ArtWork and RotoWire. Planner coverage (Gen. Plan) is 100% on ArtWork and RotoWire, indicating reliable task decomposition across domains.